ID |
Date |
Author |
Type |
Category |
Subject |
5896
|
Tue Nov 15 15:56:23 2011 |
jamie | Update | CDS | dataviewer doesn't run |
Quote: |
Dataviewer is not able to access to fb somehow.
I restarted daqd on fb but it didn't help.
Also the status screen is showing a blank while form in all the realtime model. Something bad is happening.
|
So something very strange was happening to the framebuilder (fb). I logged on the fb and found this being spewed to the logs once a second:
[Tue Nov 15 15:28:51 2011] going down on signal 11
sh: /bin/gcore: No such file or directory
[Tue Nov 15 15:28:51 2011] going down on signal 11
sh: /bin/gcore: No such file or directory
[Tue Nov 15 15:28:51 2011] going down on signal 11
sh: /bin/gcore: No such file or directory
[Tue Nov 15 15:28:51 2011] going down on signal 11
sh: /bin/gcore: No such file or directory
[Tue Nov 15 15:28:51 2011] going down on signal 11
sh: /bin/gcore: No such file or directory
[Tue Nov 15 15:28:51 2011] going down on signal 11
sh: /bin/gcore: No such file or directory
[Tue Nov 15 15:28:51 2011] going down on signal 11
sh: /bin/gcore: No such file or directory
[Tue Nov 15 15:28:51 2011] going down on signal 11
sh: /bin/gcore: No such file or directory
[Tue Nov 15 15:28:51 2011] going down on signal 11
sh: /bin/gcore: No such file or directory
Apparently /bin/gcore was trying to be called by some daqd subprocess or thread, and was failing since that file doesn't exist. This apparently started at around 5:52 AM last night:
[Tue Nov 15 05:46:52 2011] main profiler warning: 1 empty blocks in the buffer
[Tue Nov 15 05:46:53 2011] main profiler warning: 0 empty blocks in the buffer
[Tue Nov 15 05:46:54 2011] main profiler warning: 0 empty blocks in the buffer
[Tue Nov 15 05:46:55 2011] main profiler warning: 0 empty blocks in the buffer
[Tue Nov 15 05:46:56 2011] main profiler warning: 0 empty blocks in the buffer
...
[Tue Nov 15 05:52:43 2011] main profiler warning: 0 empty blocks in the buffer
[Tue Nov 15 05:52:44 2011] main profiler warning: 0 empty blocks in the buffer
[Tue Nov 15 05:52:45 2011] main profiler warning: 0 empty blocks in the buffer
GPS time jumped from 1005400026 to 1005400379
[Tue Nov 15 05:52:46 2011] going down on signal 11
sh: /bin/gcore: No such file or directory
[Tue Nov 15 05:52:46 2011] going down on signal 11
sh: /bin/gcore: No such file or directory
The gcore I believe it's looking for is a debugging tool that is able to retrieve images of running processes. I'm guessing that something caused something int the fb to eat crap, and it was stuck trying to debug itself. I can't tell what exactly happend, though. I'll ping the CDS guys about it. The daqd process was continuing to run, but it was not responding to anything, which is why it could not be restarted via the normal means, and maybe why the various FB0_*_STATUS channels were seemingly dead.
I manually killed the daqd process, and monit seemed to bring up a new process with no problem. I'll keep an eye on it. |
5895
|
Tue Nov 15 15:16:04 2011 |
kiwamu | Update | CDS | dataviewer doesn't run |
Dataviewer is not able to access to fb somehow.
I restarted daqd on fb but it didn't help.
Also the status screen is showing a blank white form in all the realtime model. Something bad is happening.

JAMIEEEE !!!! |
5894
|
Tue Nov 15 12:25:38 2011 |
kiwamu | Update | Green Locking | Y arm ALS : beat-note free run fluctuation |
Locking activity last night :
The free run beat-note in 532 nm has been measured.
However I couldn't close the ALS loop somehow.
Every time I tried closing the loop it broke the Y end PDH lock in a couple of minutes.

(Things to be done)
1. Optimization of the Y end PDH servo loop
1.1 Measurement of the arm fluctuation => to allow re-designing the servo shape
1.2 Preparation of PDH box, and temporary SR560 servo
1.3 Sanity checks on the modulation depth, reflectivity, PD dark noise and etc.,
1.4 Make the servo more robust
1.5 Some modifications on the medm screens
1.6 Activation of the temperature feedback through the realtime digital control
2. Refinement of the broadband RFPD setup
2.1 Investigation of the peak source => there was a relatively big peak around 50 MHz or so.
2.2 Noise characterization of the frequency detection system
2.3 Nicer routing of some cables.
2.4 Make two-more ADC channel connectors
2.5 Power budget on the PSL beat-note setup => estimate the expected RF level of the beat-note
2.6 Realignment of the PSL doubling and resetting of the doubling oven temperature
3. Noise budgeting
3.1 IR locked condition => measure the noise in the green beat-note system.
3.2 ALS engaged condition
3.2.0 shot noise
3.2.1 ADC noise
3.2.2 PD dark noise
3.2.3 freq. discriminator noise
3.3.4 DAC noise through the coil-magnet actuators
3.3.5 End laser suppression
3.3.6 Intensity noise
3.3.7 Thermo-elastic noise
3.3.8 Thermo-refractive noise
|
5893
|
Tue Nov 15 09:51:04 2011 |
Zach | Update | Green Locking | Y end PDH lock : UGF at 17 kHz |
Also the servo shape formed by Newfocus LB1005 looks too simple : we should have a more sophisticated servo filter (i.e. PDH box!!).
|
As promised, I will get on this this week. |
5892
|
Tue Nov 15 01:44:36 2011 |
Suresh | Update | IOO | MC WFS Servo: Open loop gain |
Quote: |
Somehow, I generically don't like the idea of lead filters for the WFS loops. We don't really need so much bandwidth. I think you should include with the servo measurements, a servo model ( on the same plot ) that matches the loop shape.
For example, this means including the 28 Hz ELP in the MC1/3 hardware and MC2 ASCPIT/YAW digital filter banks. BY comparing the model v. measurement we can determine if the cross-coupling due to imperfect output matrix is very serious or not.
In the measurements, the loop with the most low frequency gain looks the most promising.
|
WFS1_PIT servo replotted with foton data overlaid:
I included the following filters in foton:
1) Integrator: zpk([0.8],[0],0.8,"n")
2) zpk([0.8],[100,100],1,"n")
3) zpk([1:10],[3,30],1,"n")
4) ELP28
I have unwound the phase by adding or subtracting 180 to portions of the phase data.
And here is the plot for WFS1_PIT. I will repeat this process for the other three WFS loops tomorrow.

|
5891
|
Tue Nov 15 00:00:15 2011 |
Suresh | Update | IOO | MC was realigned to remove beam clipping and to accommodate PZT1 range |
[Kiwamu, Suresh]
The MC was realigned to readjust the input beam direction in pitch such that the clipping of the beam at the PSL table reduced and the railing of the PZT1 is avoided.
The current spot positions are given below on the last row:
Date |
#### |
MC1P |
MC2P |
MC3P |
MC1Y |
MC2Y |
MC3Y |
03Nov2011 |
|
0.1354 |
-0.2522 |
-0.1383 |
-1.0893 |
0.7122 |
-1.5587 |
04Nov2011 |
|
4.0411 |
4.4994 |
3.5564 |
-1.4170 |
-0.2606 |
-1.7109 |
08Nov2011 |
|
4.7341 |
4.8794 |
4.3907 |
1.3542 |
-3.0508 |
-1.7167 |
10Nov2011 |
1 |
3.9944 |
3.7676 |
6.1001 |
-1.3058 |
-3.8087 |
-1.6418 |
11Nov2011 |
1 |
3.8542 |
3.6831 |
3.0418 |
-0.8383 |
0.1550 |
-2.3841 |
11Nov2011 |
2 |
3.6876 |
2.7429 |
2.7830 |
-1.6250 |
-0.0386 |
-1.6346 |
14Nov2011 |
1 |
5.9412 |
2.7658 |
5.4806 |
-4.7676 |
0.7778 |
2.2053 |
We have quite a lot of decentering in the MC which we must try to remove by parallel transporting the beam in Pitch and Yaw..
At the current settings we might be clipping on the Faraday Isolator as we had estimated that we can allow atmost a 2mm offset in spot positions due to this constraint.
|
5890
|
Mon Nov 14 22:56:31 2011 |
kiwamu | Update | Green Locking | Y end PDH lock : UGF at 17 kHz |
[Tomotada / Kiwamu]
The open loop transfer function of the Y end PDH loop was remeasured : the UGF was found to be at 17 kHz.
The phase margin at the UGF was about 27 deg.

While the measurement we noticed that the modulation onto the laser PZT was too big
and it was creating a big AM on the reflected light with an amplitude of a few mV.
So we put a 20 dB attenuator to decrease the modulations and the reflected light became much quitter.
Also the servo shape formed by Newfocus LB1005 looks too simple : we should have a more sophisticated servo filter (i.e. PDH box!!). |
5889
|
Mon Nov 14 21:22:48 2011 |
rana | Configuration | Computers | primetime RSYNC slowing down NODUS |
nodus:elog>w; who ; date
9:20pm up 44 day(s), 5:14, 5 users, load average: 0.29, 1.04, 1.35
User tty login@ idle JCPU PCPU what
controls pts/1 9:18pm 5 -tcsh
controls pts/2 2:37pm 6:39 25:02 25:02 /opt/rsync/bin/rsync -avW /cvs/c
controls pts/3 9:14pm w
controls pts/4 4:20pm 1:56 5:02 5:02 ssh -X rosalba
controls pts/8 8:23pm 47 4:03 -tcsh
controls pts/1 Nov 14 21:18 (pianosa.martian)
controls pts/2 Nov 14 14:37 (ldas-cit.ligo.caltech.edu)
controls pts/3 Nov 14 21:14 (rosalba)
controls pts/4 Nov 14 16:20 (192.168.113.128)
controls pts/8 Nov 14 20:23 (gwave-103.ligo.caltech.edu)
Mon Nov 14 21:20:48 PST 2011
we will ask the man to stop running backups at this time of night... |
5888
|
Mon Nov 14 17:01:14 2011 |
kiwamu | Update | Green Locking | ALS feedback on MC2 |
Leaving a note on the ALS feedback before I forget:
The MC2 suspension needs to have an input for the ALS feedback in the realtime model like ETMs. |
5887
|
Mon Nov 14 15:46:14 2011 |
steve | Update | VAC | TP2's forepump changed |
The foreline pressure of TP2 was 1.4 Torr this morning. This drypump worked well for ten months.
Recently rebuilt drypump with new seal was swapped in.
This is how you do it: close V1, V4 and turn off TP2. Replace drypump and start up TP2
Set pump speed to 50 K rpm and open V4 to TP1 Note that the Maglev was not turned off because V4 was closed off only 5-10 minutes.
Open V1 the status is Vac Normal.
TP2 is rotating at 50K rpm, current pick up 0.2A, the temp is 26C and its foreline pressure 33 mTorr |
Attachment 1: drypump.png
|
|
5886
|
Mon Nov 14 12:16:41 2011 |
Jenne | Update | Computers | OAF model died for unknown reason |
I am meditating on the OAF, and had it running and calculating things. I had the outputs disabled so I could take reference traces in DTT, but the Adapt block was calculating for MCL. At some point, all the numbers froze, and the CPU meter had gone up to ~256ms. Usually it's around ~70 or so for the configuration I had (2 witness sensors and one degree of freedom enabled....no c-code calculations on any other signals). The "alive" heartbeat was also frozen.
I ssh'ed into c1lsc, ran ./startc1oaf in the scripts directory, and it came back just fine.
Anyhow, I don't know why it got funny, but I wanted to record the event for posterity. I'm back to OAFing now. |
5885
|
Mon Nov 14 11:32:02 2011 |
kiwamu | Summary | General | Goal this week |
Goal of this week : Noise budgeting on the Y arm ALS
Minimum success : bring the Y arm to the resonance by using ALS NOISE BUDGETING!!!
=> as a preparation the incident beam pointing needs to be fixed by steering the MC suspensions.
|
5884
|
Sat Nov 12 08:09:47 2011 |
rana | Update | IOO | MC WFS Servo: Open loop gain |
Somehow, I generically don't like the idea of lead filters for the WFS loops. We don't really need so much bandwidth. I think you should include with the servo measurements, a servo model ( on the same plot ) that matches the loop shape.
For example, this means including the 28 Hz ELP in the MC1/3 hardware and MC2 ASCPIT/YAW digital filter banks. BY comparing the model v. measurement we can determine if the cross-coupling due to imperfect output matrix is very serious or not.
In the measurements, the loop with the most low frequency gain looks the most promising. |
5883
|
Sat Nov 12 03:46:55 2011 |
Suresh | Update | IOO | MC WFS Servo: Open loop gain |
[Mirko, Suresh]
I closed the WFS loops and measured the transfer function from IN2 to IN1 testpoints on the WFS1_PIT filterbank.
We looked at the filter shape consisting of
1) Integrator: zpk([0.8],[0],0.8,"n")
2) zpk([0.8],[100,100],1,"n")
3) zpk([1:10],[3,30],1,"n")
The combined filter shape (along with an added pendulum filter, zpk([ ],0.8,1,"n") ) is given below

The OL Transfer function measured for WFS1_PIT loop is

The blue reference is a measurement without the third "45 deg" filter in the list above. Without it the UGF is around 1.5Hz and increasing the gain results in additional noise from the servo bump seen in the earlier elog . With it the UGF is around 3Hz.
The supression of the error signal is shown here

The other WFS loops are expected to have a similar behaviour with the exception of the MC2 QPD channels. I will measure their OLTF shortly and then proceed with the inclusion of the QPD sensors into the WFS system.
|
5882
|
Sat Nov 12 02:46:13 2011 |
Den | Update | Adaptive Filtering | stacks and ground |
We measured the coherence between the seismometer near the MC2 stack and accelerometers on the vacuum tank where MC2 is. Because accelerometers produce small signals at low frequencies, which are comparable with adc noise, we amplified the accelerometer signal by a factor of 20. We could not do it more because though adc has 40 V range, the black box that follows the channel sockets can transmit only 2.5 V max amplitude signal. Probably, this was done because old adc accepted 2 V max amplitude.


We were able to found some coherence at 0.1-1 Hz though the accelerometer signal is rather noisy. So to consider stack as a noise source is still possible. This measurement should be better done with two seismometers, one on the floor, the other on the stack. From the figure we can also see that tilt affects the x and y seismometer signals from 0.1 Hz. Green line (z-component) is much lower that red and blue lines (x and y). Tilt affects on horisontal axes of the seismometer much more than on vertical.
What we also think about is that at low frequencies mirrors start to move approximately the same and seismometers can help us to figure out small reletive displacement of the mirrors which form the MC length. We can estimate the critical frequency by presenting the ground motion as interference of surface waves with different velocities and amplitudes. For only 1 wave we have for the relation of MC length to the seismometer read out ~sin (2*pi*f/v*L). f - frequency of the wave, v -speed, L - length between the mirrors. We can see that below 1 Hz we have ~sin (f/2). At this point seismometer signal could lose coherence with MC length signal. We could try to subtract seismometer signals from corresponding axes, but gur1 and sts1 has different calibrations. Moreover, the noise floor of the seismometers might not allow us to measure the differential signal. We'll try to simulate this scenario and find seismometer calibration or measure it. We are basicly interested only in the ratio of calibraion fucntions of 2 seismometers. |
5881
|
Sat Nov 12 02:44:18 2011 |
Suresh | Update | Computers | C1IOO front end suddenly froze. Was restarted remotely |
Quote: |
[Koji Suresh]
No one was messing with the c1ioo or any other machine. The medm screens for WFS and MC alignment froze while I was working on Rossa.
There were number of red lights pertaining to c1ioo machine on the CDS_FR_STATUS screen. So we logged into c1ioo from Rossa and restarted it with 'sudo shutdown -r now'. It came back up but the C1IOO_MC_TRANS_SUM, P and Y signals were not available on the C1IOO LOCKMC screen.
I saw several messages similar to the one here
Sat Nov 12 02:09:14 PST 2011
medmCAExceptionHandlerCb: Channel Access Exception:
Channel Name: Unavailable
Native Type: Unavailable
Native Count: 0
Access: Unavailable
IOC: Unavailable
Message: Virtual circuit disconnect
Context: c1ioo.martian:43553
Requested Type: TYPENOTCONN
Requested Count: 0
Source File: ../cac.cpp
Line number: 1126
The MC autolocker script wasnt running. The heartbeat bit was not blinking on the MC_LOCKMC screen. So we manually restarted the script. Hopefully it will return to normal operation.
I restarted the fb at Sat Nov 12 02:12:19 PST 2011 in an attempt to see this resolves the problem.
It didnt.
|
The problem was resolved after I burtrestored (c1mcs c1ioo and c1rfm) epics snapshots.
|
5880
|
Sat Nov 12 02:27:00 2011 |
Suresh | Update | Computers | C1IOO front end suddenly froze. Was restarted remotely |
[Koji Suresh]
No one was messing with the c1ioo or any other machine. The medm screens for WFS and MC alignment froze while I was working on Rossa.
There were number of red lights pertaining to c1ioo machine on the CDS_FR_STATUS screen. So we logged into c1ioo from Rossa and restarted it with 'sudo shutdown -r now'. It came back up but the C1IOO_MC_TRANS_SUM, P and Y signals were not available on the C1IOO LOCKMC screen.
I saw several messages similar to the one here
Sat Nov 12 02:09:14 PST 2011
medmCAExceptionHandlerCb: Channel Access Exception:
Channel Name: Unavailable
Native Type: Unavailable
Native Count: 0
Access: Unavailable
IOC: Unavailable
Message: Virtual circuit disconnect
Context: c1ioo.martian:43553
Requested Type: TYPENOTCONN
Requested Count: 0
Source File: ../cac.cpp
Line number: 1126
The MC autolocker script wasnt running. The heartbeat bit was not blinking on the MC_LOCKMC screen. So we manually restarted the script. Hopefully it will return to normal operation.
I restarted the fb at Sat Nov 12 02:12:19 PST 2011 in an attempt to see this resolves the problem.
It didnt.
|
5879
|
Sat Nov 12 02:00:36 2011 |
Mirko | Update | Adaptive Filtering | MC-F and other signals |
Regarding http://nodus.ligo.caltech.edu:8080/40m/5867 and http://nodus.ligo.caltech.edu:8080/40m/5869 :
MC_F signal:
The measurements on p. 5867 were done using the ADC attached to the PEM computer. There was a big difference between the MC_F signal recorded directly after the server board and the signal just before the FSS board as recorded by a PEM channel.
To understand how this happens we measured the signal at different places with a spec. analyzer:
1. WIth a locked MC measuring the signal just before the PEM ADCs (meaning after a 60ft BNC cable)
2. Same position, but unlocked and seemingly dark MC
3. Locked MC, signal just before the FSS box
4. MC_F signal that is usually going into the Pentek Generic board and is recorded in C1:IOO

=> The 60ft BNC cable adds a considerable amount of noise, but doesn't fundamentally change the signal. It is weird that the signal is white from approx. 4Hz on.
Due to Jenne's measurement ( http://nodus.ligo.caltech.edu:8080/40m/5848 ) we know the TF from MCL through PD, mixer Pentek and into C1:IOO looks like this:

This is with the double HP from 15Hz on that should be in the Pentek. So one might expect a less white signal going into the FSS board...
PEM ADCs
The dark noise in the PEM ADCs is actually a factor 10 higher than in the IOO ADCs. Still ok wrt the the seismometers.
We also tried to measure essentially the dark noise of the whole seismometer readout (seismometer box, then ADC). That seemed ok, but is of limited value since the seismometer electronics behave a bit strange when there is no seismometer connected.

|
Attachment 3: Compare_signals_at_all_places.fig
|
Attachment 5: Channels_attached_to_the_PEM_ADC.fig
|
5878
|
Fri Nov 11 22:07:43 2011 |
Suresh | Update | IOO | Tried to recover the MC alignment of 4th Nov: partial success, PSL beam clipping |
I have recovered the yaw values pretty much . As the PZT1 rails in this direction perhaps this is the more relevant of the two alignments. The beam is translated in the vertical direction, but this can be easily corrected by changing the pitch of MC2
However note that if the WFS are switched on .. MC is going to follow the PSL beam.
Date |
#### |
MC1P |
MC2P |
MC3P |
MC1Y |
MC2Y |
MC3Y |
03Nov2011 |
|
0.1354 |
-0.2522 |
-0.1383 |
-1.0893 |
0.7122 |
-1.5587 |
04Nov2011 |
|
4.0411 |
4.4994 |
3.5564 |
-1.4170 |
-0.2606 |
-1.7109 |
08Nov2011 |
|
4.7341 |
4.8794 |
4.3907 |
1.3542 |
-3.0508 |
-1.7167 |
10Nov2011 |
1 |
3.9944 |
3.7676 |
6.1001 |
-1.3058 |
-3.8087 |
-1.6418 |
11Nov2011 |
1 |
3.8542 |
3.6831 |
3.0418 |
-0.8383 |
0.1550 |
-2.3841 |
11Nov2011 |
2 |
3.6876 |
2.7429 |
2.7830 |
-1.6250 |
-0.0386 |
-1.6346 |
|
5877
|
Fri Nov 11 21:09:30 2011 |
Suresh | Update | SUS | MC2 is being a little wild...WFS to blame |
Quote: |
Mirko and Den are measuring MC_F, which they will report about later, but I noticed that MC2 is totally crazy right now. It shouldn't matter that they are doing things (like unplugging the feedback to the PSL's PZT), because we actuate on the laser, not on the MC. I disabled the MC autolocker before they started working.
Anyhow, somehow MC2 got kicked up (whatever, that happens), but it won't re-damp. I think it's the WFS. The yaw output from the WFS is truely crazy.
I have disabled the WFS output / ASC input on the MC SUS screens, and MC2 was then able to damp. My disabling only the MC2 WFS input at first kicked up MC1 and 3, so I disabled all of the WFS stuff, and all 3 MC mirrors are again happy.
SURESH: FIX ME! (signed, The WFS)

|
The Problem
Turning off the WFS servo loops usually should be done using the 'mcwfsoff' script. The script takes care of switching off the integrators and Clears the History.
'mcdown' and 'mcup' scripts run the 'mcwfsoff' and 'mcwfson' scripts so when the MC unlocks the WFS servos are shutdown and restarted properly. However if the MC autolocker script is suspended by pressing the Enable/Disable switch in the LOCKMC screen and then the MC unlocks, it results in the WFS servo integrators accumulating large values. If these values are passed through the ASC filter banks the optic will get a pretty huge kick.
The Solution
I have added some indicators which will let us know if the WFS Servo Filter outputs are larger than +/-1000. When engaging the WFS loops the user has to take care to Clear History in the servo filter banks if these indicators are steadily Red. before engaging the WFS Servo loops ensure that the servo filter outputs are zeroes.
Koji and I discussed whether it would be useful to run the 'mcwfsoff' script when the Disable button is pressed in the autolocker. His recommendation is that we should keep the autolocker script simple and that user has to be cautious when switching on the WFS servos and when directing the ASC outputs to the suspensions.

|
5875
|
Fri Nov 11 14:55:47 2011 |
kiwamu | Summary | LSC | Characterization of the Power Recycled Michelson : take 2 |
Quote from #5851 |
The recycling gain is determined by the optical configuration and the optical loss in the cavity.
How much is the actual recycling gain? And how does it affect the signal extraction?
|
As Koji pointed out I made a wrong definition on the recycling gain of PRMI (Power-Recycled Michelson Interferomter).
In the correct definition the estimated recycling gain is 15.
In order to answer Koji's second question,which is about the effect on the signal extraction,
I need to scratch my head for a while.
( Give me some time..)
The value what I called "Recycling gain" must have been called "measured power build up" or something like that.
For clarity I put the definitions of the quantities.
Recycling gain :
Reflectivity of PRMI (measured by REFLDC): 
Power build up (measured by POY DC) : 
Mode Matching (MM) efficiency : 
Loss in the PRMI cavity : 
(Results of Measurement and Estimation)
Estimated recycling gain = 15
Estimated MM efficiency = 47.4%
Estimated Loss = 5.3%
Measured power build Up = 7
Measured reflectivity of PRMI = 0.5 |
5874
|
Fri Nov 11 13:35:19 2011 |
Katrin | Update | Green Locking | Feedback to ETMY |
[Kiwamu, Katrin]

Red and blue curves: frequency fluctuation of the beat node between PSL and YARM laser.
Green and broen curves: Actuation on ETMY. In ALS_CONTROL.adl ETMY filter bank 4 and 5 were switched on. Gain was 0.3
Nice reduction of the frequency fluctuation.
Y axis is in volts^2 per counts. In order to go to MHz/sqrt(Hz) you have to take the square root and then times [20Volts/(2^16)counts]*[10Hz/0.04V].
Started to scan the cavity, but this didn't work. Green light all out of lock. IR beam was badly aligned to cavity. Now, my time is over and I have to leave you.
Thanks, for your help and the nice time. |
5873
|
Fri Nov 11 13:26:24 2011 |
jamie | Update | SUS | Musings on SUS dewhitening, and MC ELP28's |
Quote: |
For the whitening on the OSEM sensor input, FM1 is linked to the Contec binary I/O. FM1 is the inverse whitening filter. Turn it on, and the analog whitening is on (bit in the binary I/O screen turns red). Turn it off, and the analog whitening is bypassed (bit in the binary I/O screen turns gray). Good. Makes sense. Either way, the net transfer function is flat.
The dewhitening is not so simple. In FM9 of the Coil Output filter bank, we have "SimDW", and in FM10, we have "InvDW". Clicking SimDW on makes the bit in the binary I/O screen gray (off?), while clicking it off makes it red (on?). Clicking InvDW does nothing to the I/O bits. So. I think that for dewhitening, the InvDW is always supposed to be on, and you either have Simulated DW, or analog DW enabled, so that either way your transfer function is flat. Fine. I don't know why we don't just tie the analog to the InvDW filter module, and delete the SimDW, but I'm sure there's a reason.
|
The input/whitening filters used to be in a similarly confused state as the output filters, but they have been corrected. There might have been a reason for this setup in the past, but it's not what we should be doing now. The output filters all need to be fixed. We just haven't gotten to it yet.
As with the inputs, all output filters should be set up so that the full output transfer function is always flat, no matter what state it's in. The digital anti-dewhitening ("InvDW") and analog dewhitening should always be engaged and disengaged simultaneously. The "SimDW" should just be removed.
This is on my list of things to do. |
5872
|
Fri Nov 11 12:32:45 2011 |
Katrin | Update | Green Locking | beat PSL - YARM laser |
[Suresh, Katrin]
Measured frequency fluctuation of the beat between PSL and YARM lasers.

Yesterday, it was very tricky to adjust the voltage offset to the slow YARM laser input to achieve the appropriate beat frequency. Today, it was much easier. During measurement beat around 25 MHz. Calibration factor 40 mV per 10 MHz. |
5871
|
Fri Nov 11 10:30:27 2011 |
rana | Update | Adaptive Filtering | MC_F |
There should be a whitening filter in the Pentek Generic DAQ board (Eurocard with 8 differential LEMO inputs). It used to be that the MC_L channel came in through here and I believe it has 2 stages of 150:15 pole:zero filters.
I don't remember if it is one or two stages, but this should be easy to measure with a function generator or by driving this input using the MC2 UL Coil monitor and doing the transfer function in DTT (as Koji and Jenne did for the demod boards). |
5870
|
Fri Nov 11 00:58:19 2011 |
Den | Update | elog | restarted elog |
Elog suspended 2 times for 1 hour. Too high frequency.
Restarted. |
5869
|
Fri Nov 11 00:55:53 2011 |
Den | Update | Adaptive Filtering | MC_F |
[Mirko, Den]
Not satisfactory work of adaptive filtering make us to think about the signals that we use. Now we try to deal with mode cleaner and analize its length. We take MC_F channel. We know that MC_F is used as a feedback signal to the laser frequency and laser changes it's frequency linear to the input modulation signal up to ~1kHz. Than is MC_F is length of MC, not velocity or acceleration. If so, it's form due to seismic noise + company of other noises + stacks and wires should be approximately like the left plot. Instead we see the right plot.
 
Possibly, left-plot form signal is not possible to transmit through the wires and adc. Most signal at medium and high frequencies would be lost because of wire and adc noise. For that reason mode cleaner length signal might be amplified at frequecnies >~20 Hz by some bandpass filter.
Where is this highpass filter and what is the form of this filter?
It might be just after the photodetector in order not to transmit real mode cleaner length through the wires. But if wires and not very noisy, it could be somewhere before ADC.
But anyway, for the laser frequency feedback the corresponding low pass filter should be used.
Where is this lowpass filter and what is the form of the filter?
We followed the mode cleaner length signal up to TT FSS and measured the mode cleaner length, that is used as an input to TT FSS. As shown http://nodus.ligo.caltech.edu:8080/40m/5867 MC_F is different from the signal that is given to TT FSS. This is not clear because we do not have smth that could effect on the signal that much before branch node and recording of MC_F. The main difference is the cut off at the MC_F signal at 3 Hz. It might be a digital filter but we do not see any filters between adc_0_0 up to MC_F test point - straight line. This means that we have an analog filter somewhere between that blue box where the branch point is and ADC. We need to find it. But at least, we do not have a lowpass filter before FSS. So it is probably after it.
So, we need to find the 3 filters that we think affect on the MC_F channel in order to figure out why we have such a bad coherence between seismic signal and mode cleaner length. |
5868
|
Fri Nov 11 00:18:53 2011 |
Zach | Update | Electronics | Precision temperature controller |
I have made a first draft of the precision temperature controller circuit, which could find use at the 40m for stabilizing EOM RFAM as well as in the Bridge labs. Please read the entry on the ATF Lab elog and give me your feedback. |
5867
|
Thu Nov 10 22:00:38 2011 |
Mirko | Update | Adaptive Filtering | Looking into MC_F & PSL misalignment |
Quote: |
[Den, Mirko]
While doing the things below we accidentally misaligned the PSL laser. Thanks to Suresh and Jenne for realigning!!
There are a lot of strange features in MC_F (see for example http://nodus.ligo.caltech.edu:8080/40m/5738 )
To get some better understanding of the signals in the control loop we looked some more into what happens to the MC feedback signal after it exits the MC servo board (D040180 see DCC).

The MC_F signal is actually the servo signal: http://nodus.ligo.caltech.edu:8080/40m/5695
The Thorlabs temperature controller is actually used in the PZT path!
We measured the LP filter in the PZT path (that is kind of mislabeled as temp.)

|
We looked into the signal from the MC servo board at different position at the PSL table.
We looked into the FB going into the temp. and PZT parts of the FB.
Temp.:

PZT:

We also looked at the signal in just in front of the FSS box No idea why the elog doesn't preview these pdfs.




Lots of extra noise there. We will check out where that comes from.
|
5866
|
Thu Nov 10 20:20:57 2011 |
Suresh | Update | IOO | MC Spot positions have shifted after accelerometer installation on MC2 chamber |
[ Jenne, Suresh ]
We were tying the fix the WFS and noticed that the PSL --> MC alignment was poor. The PMC output was also at about 0.5 instead of its optimal 0.86 . So Jenne started by first realinging the PMC input and pushed the PMC ouput to about 0.8
Then we decided to fix the PSL--> MC alignment by using the zigzag. After Jenne finished that, we realised that it was probably not the best thing to do since the MC2 might have shifted after the accelerometer installation on the MC2 chamber.
So I measured the spot positions and find that the MC2Y has shifted by about 3.6mm and MC2P has shifted by about a mm. There is also a shift of 2mm in MC3P, but hopefully it will go away when we adjust the MC2
|
|
MC1P |
MC2P |
MC3P |
MC1Y |
MC2Y |
MC3Y |
03Nov2011 |
|
0.1354 |
-0.2522 |
-0.1383 |
-1.0893 |
0.7122 |
-1.5587 |
04Nov2011 |
|
4.0411 |
4.4994 |
3.5564 |
-1.4170 |
-0.2606 |
-1.7109 |
08Nov2011 |
|
4.7341 |
4.8794 |
4.3907 |
1.3542 |
-3.0508 |
-1.7167 |
10Nov2011 |
........ |
3.9944 |
3.7676 |
6.1001 |
-1.3058 |
-3.8087 |
-1.6418 |
I am going to adjust the MC2 to recover its nominal position as marked above in green |
5865
|
Thu Nov 10 19:41:24 2011 |
Jenne | Update | SUS | Musings on SUS dewhitening, and MC ELP28's |
The following will be a stream-of-consciousness, approximately chronological story of my last hour or so of looking at screens....
In the old OAF days, we used to bypass the analog dewhitening in the coil driver path, using the XYCOMS. See, ex. elog 2548.
I began to wonder if we needed to do the same thing now. I checked several optics, to see how the switching works.
For the whitening on the OSEM sensor input, FM1 is linked to the Contec binary I/O. FM1 is the inverse whitening filter. Turn it on, and the analog whitening is on (bit in the binary I/O screen turns red). Turn it off, and the analog whitening is bypassed (bit in the binary I/O screen turns gray). Good. Makes sense. Either way, the net transfer function is flat.
The dewhitening is not so simple. In FM9 of the Coil Output filter bank, we have "SimDW", and in FM10, we have "InvDW". Clicking SimDW on makes the bit in the binary I/O screen gray (off?), while clicking it off makes it red (on?). Clicking InvDW does nothing to the I/O bits. So. I think that for dewhitening, the InvDW is always supposed to be on, and you either have Simulated DW, or analog DW enabled, so that either way your transfer function is flat. Fine. I don't know why we don't just tie the analog to the InvDW filter module, and delete the SimDW, but I'm sure there's a reason.
All optics have this setup, except MC1 and MC3. They don't have the SimDW or InvDW filter modules. Instead, in FM9 (which on all the other suspensions is SimDW, and controls the binary I/O) there is a 28Hz Elliptic Low Pass filter. The only thing I can find about these is elog 1405 where Rana talks about implementing ELP28's in MC2. But right now there is no ELP in the MC2 coil output filters. So, if Rana's old elog is to be believed, we need to fix up the ELP28 situation. But that elog was from a long time ago, so maybe things are different now? If MC1 and MC3 need the SimDW and InvDW (why wouldn't they?) then the ELP28 needs to move to another filter module. Because right now, when I click the ELP28's on and off, it changes bits in the binary I/O. Which I don't think it should. Maybe. I don't really know.
Okay. So. Now we know where everything is, and which buttons do what. Maybe not why, but at least what.
In the old world, Rob had lots and lots of trouble (elog 2027) with locking when the analog dewhitening was bypassed. But right now, I think that all of the analog dewhitening filters are bypassed, for every single optic we have. So. Which way do we want things? What's the new game plan. What's going on??
\end{stream-of-consciousness} |
5864
|
Thu Nov 10 16:44:54 2011 |
Mirko | Update | Adaptive Filtering | Looking into MC_F & PSL misalignment |
[Den, Mirko]
While doing the things below we accidentally misaligned the PSL laser. Thanks to Suresh and Jenne for realigning!!
There are a lot of strange features in MC_F (see for example http://nodus.ligo.caltech.edu:8080/40m/5738 )
To get some better understanding of the signals in the control loop we looked some more into what happens to the MC feedback signal after it exits the MC servo board (D040180 see DCC).

The MC_F signal is actually the servo signal: http://nodus.ligo.caltech.edu:8080/40m/5695
The Thorlabs temperature controller is actually used in the PZT path!
We measured the LP filter in the PZT path (that is kind of mislabeled as temp.)

|
5863
|
Thu Nov 10 16:26:46 2011 |
Mirko | Update | Computers | Firefox kills elog |
Had to restart the elog many times. For some reason firefox 8 on Win 7 kills the elog pretty consistently when trying to make a new entry. IE9 works fine .... |
5862
|
Thu Nov 10 12:28:31 2011 |
Jenne | Update | SUS | MC2 is being a little wild...WFS to blame |
Mirko and Den are measuring MC_F, which they will report about later, but I noticed that MC2 is totally crazy right now. It shouldn't matter that they are doing things (like unplugging the feedback to the PSL's PZT), because we actuate on the laser, not on the MC. I disabled the MC autolocker before they started working.
Anyhow, somehow MC2 got kicked up (whatever, that happens), but it won't re-damp. I think it's the WFS. The yaw output from the WFS is truely crazy.
I have disabled the WFS output / ASC input on the MC SUS screens, and MC2 was then able to damp. My disabling only the MC2 WFS input at first kicked up MC1 and 3, so I disabled all of the WFS stuff, and all 3 MC mirrors are again happy.
SURESH: FIX ME! (signed, The WFS)

|
5861
|
Thu Nov 10 11:52:00 2011 |
Jenne | Update | CDS | RFM signal transferring |
I am not so happy with the control signals that are coming into the OAF via the RFM/Dolphin/shmem.
The MCL/MCF signal travels via RFM from the IOO computer to the RFM model on the SUS computer, and then via dolphin to the OAF model on the LSC computer.
The MICH and PRCL signals travel via shmem from the LSC model to the OAF model, all on the LSC computer. They don't go through the RFM model.
The seismometer channels travel via shmem between the PEM model on the SUS computer and the RFM model on the SUS computer, and then via dolphin between the SUS computer and the OAF model on the LSC computer.
Each pdf shows the power spectrum and a time series of the signals in their "original" model, and in the OAF model. The seismometer is the only one that seems fine. The time series match, except for a delay which is not surprising, since the signals have to travel. The other signals seem pretty distorted. What is going on??? Why can we trust some, but not all, of the signals that move between models and between computers???
(This data was all taken while the MC was locked, but MICH and PRCL were not. I don't think this should have any effect on the signal transfer though).
The MCL isn't soooo bad, so maybe we can keep moving forward with it, but I'm concerned that we're not really going to be successful OAF-ing the other degrees of freedom if the signals are so distorted. |
Attachment 1: OAF_rfm_signals_MCL.pdf
|
|
Attachment 2: OAF_rfm_signals_MICH.pdf
|
|
Attachment 3: OAF_rfm_signals_PRCL.pdf
|
|
Attachment 4: OAF_rfm_signals_GUR1X.pdf
|
|
5860
|
Thu Nov 10 05:54:23 2011 |
kiwamu | Update | Green Locking | Beat-note detected : PSL vs Y arm |
[Katrin / Kiwamu]
The beat-note between the PSL green laser and the Y end green laser was successfully detected.
The detection was done by the new broad-band RFPD.
The next step will be an extraction of the frequency fluctuation signal using the delay-line-mixer frequency discriminator.
Here is a picture of the RF spectrum analyzer displaying the direct output signal from the broad-band RFPD.
The beat-note was moving around 100 MHz with an RF power of -36 dBm. The frequency fluctuation was about +/- 7MHz in a time scale of 1 sec or so.

(What we did)
+ Connected a BNC cable which goes from the c1iscey's DAC to the laser slow input
=> this enables a remote control of the laser frequency via the temeperature actuation
+ Realigned the beam pointing of the Y end green laser
+ Installed all the necessary optics on the PSL table
=> currently the PSL green light is adjusted to completely S-polarization
+ readjusted the mode matching telescopes
=> the Y green beam becomes the one with a long Rayleigh range
+ Health check on the broad-band RFPD to see if it is working
+ Installed the BB-RFPD with a +/-15V power supply
+ Fine alignment of the beam combining path
+ Fine tuning of the Y end laser temperature
=> T_PSL = 31.72 deg when the slow FSS feedback is zero.
=> Based on Bryan's measurement (see #elog) the Y end laser temperature was adjusted to 34.0 deg by applying an offset to the slow input.
+ Found the beat note at 100 MHz or so.
=> optimizing the alignment of the beam combining path by maximizing the peak height of the beat-note.
=> maximum peak height observed with an RF spectrum analyzer was about -36 dBm.
|
5859
|
Wed Nov 9 21:48:43 2011 |
Suresh | Update | IOO | WFS Servo included into the MC_Autolocker |
The WFS servo loop will come on 5 seconds after the MC is locked.
I have uncommented the lines in the mcup script which turn on the WFS servos. But I shifted their location to the part after the MC is locked. |
5858
|
Wed Nov 9 21:32:38 2011 |
Mirko | Update | Adaptive Filtering | Put accelerometers 4-6 on top of MC2 tank |
Put the accelerometers on top of MC2. Orientated as the arms,GUR1 and STS1:

Should be fixed a bit more rigidly.
Looking into the signals at a quiet time:

Hmm. Either the acc. are mislabeled or there is really bad x-y coupling. The connectors in the back of the acc. power supply / amplifier box are in ascending order. |
Attachment 3: Coherence_quiet_time.fig
|
5857
|
Wed Nov 9 21:21:30 2011 |
Suresh | Update | IOO | MC WFS: Output matrix determined with loops closed. |
With the loops closed I ran the $SCRIPTS/MC/WFS/senseWFSoutMATRX script and analysed the lockin outputs with $SCRIPTS/MC/WFS/matlab/wfsmatrix3.m. I had to edit both the setupWFSlockins and the sensWFSoutMATRX scripts because in the past we used to switch on / off the ASC filter bank GAINs on the MC suspensions to start / stop the lockin excitation. We cannot do this any more since these the WFS feedback signals have to get through these filters while the WFS loops are closed. So the current, more sensible, scheme is to set the appropriate elements to 1 / 0 in the C1IOO_LKIN_OUT_MTRX.
Note: The senseMCdecenter script will also have to be ammended in the same manner.
The lockin outputs measured are (reduceddata):
wfs1P wfs2P mc2tP wfs1Y wfs2Y mc2tY
MC1P -10.3694 7.0642 -10.2133 -0.1025 0.4653 -0.0000
MC2P 8.2838 21.5141 8.4102 -0.2215 0.0734 0.0000
MC3P 9.4804 6.0835 9.6346 -0.0080 0.0366 -0.0000
MC1Y -0.7339 -1.4498 -0.6175 -11.7502 -13.0480 0.0004
MC2Y 0.9004 0.6645 1.0554 25.6083 7.3399 -0.0046
MC3Y -0.2914 2.1573 -0.1829 -2.1130 14.3038 -0.0000
After inverting and normalising a subset of the above matrix ( done in the wfsmatrix3.m ) we obtain the following output matrix coefs:
|
WFS1P |
WFS2P |
MC1P |
-1.0 |
0.82 |
MC2P |
0.15 |
1.0 |
MC3P |
0.62 |
0.02 |
|
WFS1Y |
WFS2Y |
MC1Y |
-0.11 |
-0.56 |
MC2Y |
1.00 |
-0.17 |
MC3Y |
-0.62 |
1.00 |
Apart from a negative sign (introduced by the negative gains in the WFS servo filters ) these values are quite close to the actuation vectors determined in open loop.
I have plugged these values into the WFS output matrix. Will determine the open loop gain later when there arent so many people stomping around the MC.
|
5856
|
Wed Nov 9 20:35:58 2011 |
Mirko | Update | Adaptive Filtering | Seismic noise injection into the MC |
Very elaborated measurement ;-)
On 11-11-08:
18:40 Stomp near STS1 for 2mins
18:47 Jump near GUR1 for 2mins
18:52 Walk from MC2 approx. half-way to vertex for 2mins
Tried to see if jumping / stomping the ground near STS1 / vertex or GUR1 / MC2 would show up in the seismometer or MC length data.
In GUR1 jumping / stomping clearly shows up in the timeseries. Also it clearly shows up as a low frequency signal if you walk to a position near MC2. E.g. walk from the vertex to MC2. Stop near the cones. Gives a big dip on GUR1X, that recovers in 10-20sec if you remain stationary. Big "hill" if you come from x-arm end and stop on the x side of MC2. So probably lots of tilt to GUR1X coupling at low frequencies.
Nothing was really visible in spectra (see below).
Resonances:
There appear to be a lot of resonances in the 10-20Hz range, see e.g. 1st attached pic.
Coherence:
Looking at the coherence of difference axis of the seismometers. Kind of dirty measurement, could have all kinds of reasons.
Quite a bit of coherence in STS1 at 5-6Hz. Possibly limiting the STS1X to MC-F coherence to up to 4Hz?


|
Attachment 1: Inj_spectra_at_GUR1_all_DOFs.fig
|
Attachment 2: Inj_spectra_at_GUR1_all_DOFs.png
|
|
Attachment 3: Inj_spectra_at_STS1_all_DOFs.fig
|
Attachment 4: Inj_spectra_at_STS1_all_DOFs.png
|
|
Attachment 7: Coherence_GUR.fig
|
Attachment 8: Coherence_STS1.fig
|
5855
|
Wed Nov 9 19:08:18 2011 |
Suresh | Update | elog | restarted elog |
Elog was not responding and was restarted. |
5854
|
Wed Nov 9 18:02:42 2011 |
jamie | Update | CDS | ISCEX front-end working again (for the moment...) |
The c1iscex IO chassis seems to be working again, and the iscex front-end is running again.
However, I can't say that I actually fixed the problem.
Originally I thought the timing slave board had died by the fact that the front LED indicators next to the fiber IO were out. I didn't initially consider this a power supply problem since there were other leds on the board that were lit. I finally managed to track down Rolf to give downs the OK to pull the timing boards out of a spare IO chassis for us to use. However, when I replaced the timing boards in the chassis with the new ones, they showed the exact same behavior.
I then checked the power to the timing boards, which comes off a 2-pin connector from the backplane board in the back of the IO chassis. Apparently it's supposed to be 12V, but it was only showing ~2.75V. Since it was showing the same behavior for both timing boards, I assumed that the issue was on the IO chassis backplane.
I (with the help of Todd Etzel) started pulling cards out of the IO chassis (while power cycling appropriately, of course) to see if that changed anything. After pulling out both the ADC and DAC cards, the timing system then came up fine, with full power. The weird part is that everything then stayed fine after we started plugging all the cards back in. We eventually got back to the fully assembled configuration with everything working. But, nothing was changed, other than just re-seating all the cards.
Clearly there's some sort of flaky connection on the IO chassis board. Something is prone to shorting, or something, that overloads the power supply and causes the voltage supply to the timing card to drop.
All I can do at this point is keep an eye on it and go through another round of debugging if it happens again.
If it does happen again, I ask that everyone please not touch the IO chassis and let me look at it first. I want to try to poke around before anyone giggles any cables so I can track down where the issue might be. |
5853
|
Wed Nov 9 17:41:17 2011 |
Jenne | Update | Computers | ETMX restored |
Jamie did computer magic. I burt restored scxepics, and restored ETMX damping. |
5852
|
Wed Nov 9 16:49:17 2011 |
kiwamu | Update | Green Locking | Y end laser temperature with slow input connected |
Indeed it is strange. I took a quick look at it.
In order to recover the same condition (e.g. the same amount of the reflected DC light and the same temperature readout),
it needed to have +8.9V in the slow input from the DAC through EPICS.
Obviously applying an offset in the slow input to maintain the same condition is not good.
It needs another solution to maintain the sweet frequency where the frequency of the PSL and the Y end laser is close in a range of 200 MHz.
Quote from #5797 |
Plugging in the thermal feedback BNC cable to the laser reduced the DC voltage of the green PDH photo diode from 3.12 V to 1.5V off resonance.
|
|
5851
|
Wed Nov 9 16:29:36 2011 |
Koji | Summary | LSC | Characterization of the Power Recycled Michelson |
Difficult to understand.
The mode matching does not change the recycling gain. It changes the coupling of the incident light into the cavity.
It changes the reflected and accumulated power, but the recycling gain is not affected.
The recycling gain is determined by the optical configuration and the optical loss in the cavity.
In the actual measurement, of course, we should take both of the loss and the mode matching into account.
But this is the issue of the measurement method.
The essential questions are:
How much is the actual recycling gain? And how does it affect the signal extraction? |
5850
|
Wed Nov 9 16:03:21 2011 |
kiwamu | Summary | General | Goal this week |
Goal of this week : ALS on the Y arm
Minimum success : Detection of the green beatnote between the freq-doubled PSL and the Y arm transmitted light
|
5849
|
Wed Nov 9 14:49:07 2011 |
kiwamu | Summary | LSC | Characterization of the Power Recycled Michelson |
EDIT by KI:
The definition of the recycling gain is wrong here.
See the latest entry (#5875)
Here is a summary about the Power Recycled Michelson (PRMI).
It seems the mode matching is also one of the greatest contributor on the low recycling gain.
(Estimated parameters)
+ Loss = 5.3% (or effective reflectivity of 93.28% in Michleson) => Under coupling !!
+ Mode matching efficiency = 47.4 % => Really bad !!
With these values we end up with a recycling gain of 7 and a normalized REFLDC of 0.5 as observed (#5773).
Also according the incident beam scan measurement (#5773) the loss is NOT a local effect like a clipping, it is more like uniformly distributed thing.
As for the mode matching, the number indicates that approximately the half of the incident light is coming back to the REFL port without interacting with PRMI.
This is bad because the non-mode-matched light, which is just a junk light, is entering into the photo detectors unnecessarily.
In the worst scenario, those junk light may create a funny signal, for example a signal sensitive to the alignment of PRM.
(Estimation method)
The method to estimate the loss and the MM (Mode-Matching efficiency) is essentially the same as before (#5541).
One difference from the previous estimation is that the I used more realistic parameters on the transmissivity of ITMs and PRM :
PRM : T = 5.637 % (see the 40m wiki)
ITM : T = 1.384 % (see the 40m wiki)
In addition to the basic calculations I also made plots which are handy for figuring out where we are.
Quantities we can measure are the reflected light from PRMI and the recycling gain using the REFL PD and the POY PD respectively.
So I wanted to see how the loss and MM can be estimated from the measured REFL DC and recycling gain.
The plots below are the ones.

[Loss map]
The first figure shows a contour map of the loss as a function of the measured REFL DC and recycling gain.
The white area is a place where no proper solutions can be found (for example MM can get to more than 100 or loss becomes negative).
The star mark in the plot corresponds to the place where we are now. Obviously the loss is about 5%.
If we somehow decrease the amount of the loss the star mark will mostly go up in the plot.
[MM map]
The second figure shows a contour map of the MM as a function of the measured REFL DC and recycling gain.
The X and Y axis are exactly the same as that of the first plot. Again the star mark represents the place where we are.
We are currently at MM=47%
(Solutions)
Here are some solutions to bring the recycling gain higher.
We don't work on these things immediately since it requires opening of the chambers again and it will take some times.
But we should think about those options and prepare some stuff for a coming vent.
+ Refinement of the position of the mode matching telescopes. => The Recycling gain can go up to 15.
=> Assuming the loss in the cavity doesn't change, the star mark in the first plot will go to the left hand side along the "0.05" black solid line.
=> However PRMI will be still under coupled.
=> Needs an estimation about which way we move the telescopes.
+ Locate the place of the dominant loss source and reduce it somehow.
=> The recycling gain will be more than 18 if the loss reduces by a factor of more than 5.
=> Needs a clever way to find it otherwise we have to do it in the classical way (i.e. white light and trying to find dirty surfaces) |
5848
|
Wed Nov 9 14:23:35 2011 |
Jenne | Update | Adaptive Filtering | OAF MC Delay Measurement |
As described in elog 2063 and the mevans document, we need to measure the TF of the OAF's plant, to find the delay.
At DC, the phase is 2.5deg, and at 32Hz, the delay is -4.6Hz (extrapolated from the points at ~30deg and ~38deg). The DTT file is in /users/Templates/OAF/OAF-MCL-Delay-9Nov2011.xml .
This gives a phase lag of 7.1deg at the Nyquist freq.
7.1 / 180 * 32 = 1.26, so ~1 cycle delay. Not so much. The new ADCs are waaaay faster than the old 110Bs. Yay! |
Attachment 1: OAF-MCL-Delay-9Nov2011.pdf
|
|
5847
|
Wed Nov 9 13:44:04 2011 |
Mirko | Update | Computers | NDS1 missing channels in matlab |
The Matlab NDS1 access seems to only work for some channels. With other channels it just hangs 'Busy' and does nothing.
Below you can see some channels that make matlab hang. Everyting happened on allegra. I tried compiling the NDS1 sources (from https://www.gravity.phy.syr.edu/dokuwiki/doku.php?id=ligodv:nds1_ligodv_install ) into mex files myself. Same result. I
a=NDS_GetChannels('fb:8088'); %/cvs/cds/caltech/apps/linux64/share/matlab/NDS_GetChannels.m
%data=NDS_GetData({'C1:IOO-MC_F_DQ'},1004826500,100,'fb:8088',a) %Works
%data=NDS_GetData({'C1:IOO-WFS1_PIT_IN1_DQ'},1004826500,100,'fb:8088',a) %Works
data=NDS_GetData({'C1:LSC-AS11_I_OUT'},1004826500,100,'fb:8088',a) %Doesn't work, hangs
%%%which NDS_GetData.m: /cvs/cds/caltech/apps/linux64/share/matlab/NDS_GetData.m |
5846
|
Wed Nov 9 12:05:08 2011 |
Katrin | Update | Green Locking | YARM error signal and feedback signal |
error signal = signal measured behind the low-pass filter
feedback signal = output of the gain servo, going to the PZT
First of all both signals in V/sqrt(Hz) just in case I mess up the next calibration step.

The 60 Hz line (and its multiple) are a new feature. They show up as soon as the feedback loop is closed. So far, I couldn't find their origin.
For the next calibration step:
- width of a typical error signal, i.e. the frequency band width of the carrier slope, ~1.4 kHz
- height of a typical error signal 182 mV
|